OpenStudy (jannine24):

need help solving this function y= x^2 -4 and also y=-3x^2-6x-5

OpenStudy (anonymous):

are the functions equal to zero?

OpenStudy (jannine24):

no i have to solve then graph them

OpenStudy (jannine24):

i have tried them both but i am not sure they are correct

OpenStudy (anonymous):

what are you solving for?

OpenStudy (stamp):

Since they both are in terms of y, set them equal to each other and then solve for x. This will be the x value at which the two functions intersect. Plug this x value back into either equation to find the point x, y at which the lines merge.

OpenStudy (jannine24):

i need ot find the vertex if any the x intercept and y intercept

OpenStudy (anonymous):

can you rephrase please

OpenStudy (jannine24):

i need to find the vertex the xintercept and y intercept which would make them a parable

OpenStudy (whpalmer4):

x-intercept is the point(s) where y = 0. y-intercept is the point(s) where x = 0. To find the x-intercept of \( y = x^2 - 4\), solve\[0=x^2-4\]\[x^2=4\]What values satisfy that equation? Hint: remember that a negative number times a negative number gives you a positive number.

OpenStudy (jannine24):

it would be 2

OpenStudy (jannine24):

or -2

OpenStudy (whpalmer4):

Correct, those are the x-intercepts of y = x^2 - 4. Now can you find the y-intercept(s)? Proceed similarly, except this time you set x = 0 and find the value of y that makes it happen.

OpenStudy (whpalmer4):

(it's easy!)

OpenStudy (jannine24):

y would be 0,-4 but wouldn't that be the vertex

OpenStudy (whpalmer4):

Yes, the y-intercept is at (0, -4). This is a parabola that is shifted down 4 units from the origin.

OpenStudy (jannine24):

so then their is no vertex

OpenStudy (whpalmer4):

Oh, there's always a vertex if it's a parabola! Can you tell me what the vertex of a parabola is? (not the value, just describe it)

OpenStudy (jannine24):

the vertex of parabola is point where it crosses the axis

OpenStudy (whpalmer4):

Um, no, not necessarily. Though many of the parabolas you've probably encountered might have their vertices there, it isn't generally true.

OpenStudy (whpalmer4):

 x marks the spot of the vertex of the parabola, even though it doesn't happen to fall on either axis

OpenStudy (jannine24):

so they can be anyplace then

OpenStudy (jannine24):

they show the highest or lowest point

OpenStudy (whpalmer4):

If the parabola points up or down, yes, it's the bottom or the top. But you can have a parabola that is on its side, too! \(y^2=x\) would be an example.

OpenStudy (whpalmer4):

So we better figure out the general way to find the vertex, eh? :-)

OpenStudy (jannine24):

i would use the formula -b/2a

OpenStudy (jannine24):

here their is no b so it would be -0/2(1)

OpenStudy (whpalmer4):

That would be a good way. The vertex of y = x^2 -4 is at (0,-4). We get the 0 from the formula, and the -4 from plugging the 0 into the equation of the parabola.

OpenStudy (whpalmer4):

(btw, "there is no b" is correct, "their" is a possessive pronoun, as in "that is their homework for the day")

OpenStudy (jannine24):

yes that is what i did

OpenStudy (whpalmer4):

Now how about that other parabola?

OpenStudy (whpalmer4):

\[y=-3x^2-6x-5\]

OpenStudy (whpalmer4):

We need to find x-intercept(s), y-intercept(s), and vertex. Shall we race? :-)

OpenStudy (jannine24):

this one i found the vertex

OpenStudy (jannine24):

which is -1,-2

OpenStudy (jannine24):

and yintercept is 0,-5

OpenStudy (jannine24):

then i solved with quadratic formula of -b plus or minus the square root of b^2 -4ac/2a

OpenStudy (whpalmer4):

so what do you get for the x-intercepts?

OpenStudy (whpalmer4):

(this might be regarded as a bit of a trick question)

OpenStudy (jannine24):

when i did this i ran into a problem

OpenStudy (whpalmer4):

okay...what were the solutions the quadratic formula gave you?

OpenStudy (jannine24):

you cannot take the square root of -24 so then thier is no answer

OpenStudy (whpalmer4):

Well, you can take the square root of -24. You just get what's called an imaginary number! As if the other kind are not imaginary, right? :-)

OpenStudy (whpalmer4):

I take it you haven't encountered \( i = \sqrt{-1}\) yet?

OpenStudy (jannine24):

no i have not

OpenStudy (jannine24):

i was told if we have a negative in the square root we cannot do the problem then it is done

OpenStudy (whpalmer4):

Okay, well, as it turns out, I'm reading a book called "An Imaginary Tale: The Story of \(\sqrt{-1}\)"! It seems that mathematicians of old (prior to 1600 or so) had a real fear of negative numbers, and would write their problems so that no negative numbers appeared.

OpenStudy (jannine24):

sound like a good book

OpenStudy (whpalmer4):

Then someone figured out that it was actually a pretty useful concept... Anyhow, the 10 cent tour: if you let \(i^2 = -1\rightarrow i=\sqrt{-1}\) then you can factor your negative number under the square root sign like this: \[\sqrt{-24} = \sqrt{-1*24} = \sqrt{24} * \sqrt{-1} = 2i\sqrt{6}\]

OpenStudy (whpalmer4):

And there's a really cool result involving \(i, \pi, e, -1, 0\) (the big 5 numbers in math, you might say): \[e^{i\pi} - 1 = 0\] Who would think that those numbers from all over, so to speak, would come together in such an elegant way? What does the ratio of the circumference of a circle to its diameter have to do with the base of the natural logarithm, or the square root of -1? But there it is!

OpenStudy (jannine24):

we were not taught to do this though..who wrote this bool

OpenStudy (jannine24):

sorry book

OpenStudy (whpalmer4):

Anyhow, back to our parabola. You don't get any answers that are real numbers, only answers that have \(i\) in them. That means there are no x-intercepts in this case (which is why I called it a trick question when I asked you what they are).

OpenStudy (jannine24):

so i was right then no xinterecept

OpenStudy (jannine24):

i am not crazy then

OpenStudy (whpalmer4):

Here's a graph of our parabola:

OpenStudy (jannine24):

to graph this how would i even do it

OpenStudy (jannine24):

it would go down but not touching the xaxis

OpenStudy (whpalmer4):

The book is by a guy named Paul Nahin. It's got a lot of complicated math in the first chapter, and more in the rest of the chapters, so it might be a difficult book to read, but the history parts and the descriptions of many different ways in which \(i\) comes into play is interesting.

OpenStudy (whpalmer4):

Graph the way you graph any other function: make a table of x and y values and plot the points. The trick is figuring out which area of the infinite x-axis and y-axis is likely to be interesting!

OpenStudy (jannine24):

could i use say 1 for x and plug it into the problem to add points

OpenStudy (whpalmer4):

You can sort of assume at your level that probably the interesting area isn't going to be too far from the origin, but eventually you may not be able to make that assumption. Yes, I would make a table of say x = -5, -4, -3, -2, -1, 0, 1, 2, 3, 4, 5 and then grind out the values of y for each x. Maybe do every other one at first. I always do 0, because that usually is pretty easy to evaluate :-)

OpenStudy (jannine24):

thank you